# AnnaAgent: Dynamic Evolution Agent System with Multi-Session Memory for Realistic Seeker Simulation

### v1

@@ 1 Introduction
Undeniably these works have made a significant contribution to the study of seeker simulation, addressing issues like CAs’ easily endorsing advice, sycophantic obedience Wester et al. (2024); Black (2023), and emotional flatness Wang et al. (2023a); Balcombe (2023). However, CAs designed with existing methods are still not realistic enough, blocking the further development of AI for mental health research. As shown in Figure 1(a), A real seeker is usually more sensitive due to the psychological disorders suffered, and there is usually a noticeable dynamic evolution of the emotional state during the counseling process. Nevertheless, existing simulators fail to replicate this dynamic evolution. Throughout the counseling process, their emotions are usually maintained at the initially set state, without noticeable changes. Moreover, psychological counseling is typically a long-term, multi-session task, and counselors usually need to review previous session topics, discuss recent circumstances with seekers to build trust, assess progress, and analyze in depth the core dilemmas of seekers Barkham et al. (2015). However, existing methods do not provide CAs with multi-session memories, which may lead to confusion and hallucinatory responses from CAs when counselors mention the topic of previous sessions. In Figure 1(b), when the counselor attempts to verify the effectiveness of the suggestions made in the previous counseling session, the simulator, who has no multi-session memory, gives a relevant but incorrect response. Thus, to simulate seekers more realistically, we present two pressing challenges: Dynamic Evolution and Multi-session Memory. Figure 1 visualizes these two challenges with examples of scenes from counseling.
@@ 3.2 Dynamic Evolution
Based on the previous description, dynamic evolution focuses on the dynamic control of the seeker’s emotions and chief complaints during a counseling session.
@@ Appendix E Questions for Long-term Memory
Since the probability of long-term memory triggering is relatively low in general, we designed these questions above in order to verify the role of long-term memory more effectively. The probability of triggering long-term memory was effectively increased by asking virtual seekers to focus on questions related to previous sessions.

### v2

@@ 1 Introduction
Undeniably, these works have made a significant contribution to the study of seeker simulation, addressing issues like CAs’ easily endorsing advice, sycophantic obedience Wester et al. (2024); Black (2023), and emotional flatness Wang et al. (2023a); Balcombe (2023). However, CAs designed with existing methods are still not realistic enough, blocking the further development of AI for mental health research. As shown in Figure 1(a), A real seeker is usually more sensitive due to the psychological disorders suffered, and there is usually a noticeable dynamic evolution of the emotional state during the counseling process. Nevertheless, existing simulators fail to replicate this dynamic evolution. Throughout the counseling process, their emotions are usually maintained at the initially set state, without noticeable changes. Moreover, psychological counseling is typically a long-term, multi-session task, and counselors usually need to review previous session topics, discuss recent circumstances with seekers to build trust, assess progress, and analyze in depth the core dilemmas of seekers Barkham et al. (2015). However, existing methods do not provide CAs with multi-session memories, which may lead to confusion and hallucinatory responses from CAs when counselors mention the topic of previous sessions. In Figure 1(b), when the counselor attempts to verify the effectiveness of the suggestions made in the previous counseling session, the simulator, which has no multi-session memory, gives a relevant but incorrect response. Thus, to simulate seekers more realistically, we present two pressing challenges: Dynamic Evolution and Multi-session Memory. Figure 1 visualizes these two challenges with examples of scenes from counseling.
@@ 3.2 Dynamic Evolution
Based on the previous description, dynamic evolution focuses on the dynamic control of the seeker’s emotions and chief complaints during a counseling session. We set up a reminder after each round of dialogue to remind the virtual seeker of the current state of emotion and chief complaint (see Figure 9(b)).
@@ Appendix E Questions for Long-term Memory
Recognizing that the spontaneous retrieval of long-term memory is typically an infrequent event, we strategically designed the aforementioned questions to create a more reliable method for its verification. Our approach aimed to overcome the challenge of low trigger probability by explicitly directing the virtual seekers to engage with inquiries linked to their experiences in previous sessions. This targeted methodology substantially increased the likelihood of activating long-term memory, thereby enabling a more effective and conclusive analysis of its functional role within our framework.